OpenStudy (nincompoop):

what causes the nucleus of an atom to weigh less than the sum of its isolated component particles?

OpenStudy (anonymous):

E=mc^2

sam (.sam.):

Nuclear fission

OpenStudy (nincompoop):

\[∆E = ∆m c^2\]

OpenStudy (nincompoop):

a change in energy as it becomes more stable, the mass decreases?

sam (.sam.):

The nucleus of an atom splits into smaller parts, producing free neutrons and photons. Hence \(E=mc^2\), and releasing a very large amount of energy. The masses are typically with a mass ratio of products of about 3 to 2,

OpenStudy (anonymous):

It's called the "mass defect" - see: http://en.wikipedia.org/wiki/Binding_energy

OpenStudy (nincompoop):

I know what it is called, but what causes the phenomena to occur is what I need a little bit of information on. I understand how the equation of relativity easily explains it, but it does not account that gluons are supposedly massless.

OpenStudy (nincompoop):

am I just talking nonsense trying to delve into this by missing some fundamentals?

OpenStudy (anonymous):

The mass of the gluons don't matter. They are force-carriers. What matters is the difference in energy between bound particles and unbound.

OpenStudy (nincompoop):

thanks cliff.

OpenStudy (anonymous):

Some of the fundamentals are that particles always seek the lowest potential energy configuration. In a stable nucleus, the potential energy is lower than if the particles were separate (because there is a force of attraction due to the gluons). With increased separation, the force is able to do more work and the potential energy increases. This is analogous to lifting an object against a gravitational field, the potential energy increases. For nucleons that fission from a nucleus, the increase in potential energy is seen in the form of increased mass (because mass and energy are kinda the same thing.)

OpenStudy (nincompoop):

from what I gathered, you are suggesting that potential energy has something to do with increased or decrease in mass.

OpenStudy (anonymous):

E=mc^2

OpenStudy (nincompoop):

lol thanks i knew that it'd come back to that

OpenStudy (nincompoop):

"energy"

OpenStudy (anonymous):

Consider this for extra silliness: Quarks tend to come in bound pairs or triplets. The energy of attraction between quarks is so large that if you attempt to separate two quarks, the amount of energy required will create two more quarks.

OpenStudy (nincompoop):

i learned that from Neil deGrasse's lectures.

OpenStudy (anonymous):

Nice. :-) I recommend some Lawrence Krauss too.

OpenStudy (nincompoop):

I know. I was looking at Krauss' book and his past works, but right now I am bogged down with too much stuff for me to be able to sit down and digest it fully.

OpenStudy (nincompoop):

thanks for your help.